# Circuits and Backdoors: Five Shades of the SETH

### v1

@@ 3.2 Equivalent Formulation of the SETH for Max-SAT
Take such a component $C$ and for each assignment $s$ to its $\delta$ neighbor variables in $M$, calculate the assignment to the variables of $C$ that maximizes the number of clauses involving variables of $C$ which are satisfied (this can be done by going through the $2^{\sigma}$ possible assignments to the variables of the component). Let $w_{s}$ be the number of such clauses. Our goal is to remove the variables of $C$ from $\phi$ and add some clauses involving the variables of $N(C)$ such that for each assignment $s$ the total weight of satisfied new clauses is proportional to $w_{s}$. More precisely, we do the following: for each assignment $s$ to the variables of $N(C)$ we construct a clause falsified by this assignment (and only this assignment) to $N(C)$, and give it weight $\sum_{s^{\prime}\neq s}w_{s^{\prime}}$, that is give it weight that is equal to the sum of the weights corresponding to all other assignments to $N(C)$. We now observe that if we use the assignment $s$ for $N(C)$, the total weight we obtain from the new clauses is $w_{s}+(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$, that is, the contribution of each assignment (including $s$) is counted $22^{\delta}-$ 2 times, but $w_{s}$ is counted once more, as a $w_{s}$ term appears in the weight of all satisfied clauses. We increase the target value of the Max-SAT instance by $(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$ and implement the weights by taking copies of the constructed clauses. After performing this process exhaustively we are left with a Max-SAT instance with maximum arity $\max\{k,\delta\}$ and size at most $O(2^{\sigma+\delta}|\phi|)$, so executing the supposed algorithm gives the lemma.
@@ 4 Modulators to Constant Width and Linear Depth Circuits
In this section we present a sequence of reductions which show that several seemingly distinct fine-grained complexity hypotheses are actually the same. Our results our summarized in Theorem 4.1 below.
@@
    (LD-C-SETH is false) There is an $\epsilon>0$ and an algorithm which takes as input a boolean circuit of size $s$ on $n$ inputs with depth at most $\epsilon n$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
@@
    (pwM-SETH is false) There is an $\epsilon>0$ such that for all fixed $c>0$ we have an algorithm which takes as input a CNF formula $\phi$ and a size $m$ modulator of $\phi$ to pathwidth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
@@
  5. 5.
    (W[SAT]-SETH is false) There is an $\epsilon>0$ such that there is an algorithm which takes as input a general Boolean formula $\phi$ on $n$ variables and an integer $k$ and decides if $\phi$ has a satisfying assignment that sets exactly $k$ variables to True in time $n^{(1-\varepsilon)k}|\phi|^{O(1)}$.
Among the five statements of Theorem 4.1 three have an obvious chain of implications: solving SAT parameterized by the size of a modulator to logarithmic tree-depth is at least as hard as solving it for a modulator to constant treewidth, which is at least as hard as for a modulator to constant pathwidth. These relations follow readily from the fact that $\textrm{pw}\leq tw$ and $\textrm{td}\leq\textrm{tw}\log n$ for all graphs. What is perhaps a bit surprising is that these three parameters turn out to be of equivalent complexity.
In order to add the remaining chains of implications to complete Theorem 4.1 we need several steps. The one that is conceptually the least obvious is perhaps the implication 2$\Rightarrow$1. Here we need to prove that the satisfiability problem for any circuit (with the stated restriction on its depth) can be reduced to SAT with a modulator to only constant pathwidth. The idea now is that the variables of the modulator will correspond to the inputs of the circuit, so we need to encode the workings of the circuit via a formula of constant pathwidth. This sounds counter-intuitive, since evaluating a circuit is a P-complete problem, but constant pathwidth SAT instances can be solved in logarithmic space. This observation makes it clear why it is crucial that we have placed a restriction on the depth of the circuit. In order to achieve our goal we rely on a classical result from computational complexity theory known as Barrington’s theorem [5] which allows us to transform any boolean circuit of depth $d$ into a branching program of length $2^{O(d)}$ and constant width. The fact that the branching program has constant width means that its workings can be encoded via a constant pathwidth formula, while the fact that the depth of the input circuit is at most $\epsilon n$ means that the exponential blow-up in the size can be tolerated by carefully dealing with the involved constants.
The other interesting implication is 5$\Rightarrow$4. To better explain this, suppose we were instead trying to prove 1$\Rightarrow$4. In that case we would try to construct a circuit whose inputs are the variables of the modulator given to us in statement 4. For this, we would need to somehow eliminate all the remaining variables, using the fact that they form a graph of logarithmic tree-depth. For this we use an inductive construction which proceeds bottom-up, constructing for each node of the tree a formula that eliminates all variables in its sub-tree. To do this we try both values of the variable and construct two copies of the formulas we have for its children, one for each value. This construction is of course exponential in the tree-depth, but since the tree-depth is $O(\log|\phi|)$, the result is polynomial in the input formula. To obtain instead 5$\Rightarrow$4 we use standard tricks that encode an assignment to $n$ variables via a weight $k$ assignment to $2^{n/k}$ variables.
Finally, for 1$\Rightarrow$5 we can construct a circuit with $k\log n$ inputs, encoding the $k$ variables to be set to True in the sought weight-$k$ assignment. A subtle point here is that we need to ensure that the circuit we construct does not have too large depth. However, for this we can rely on the fact that the W[SAT]-complete problem of statement 5 deals with boolean formulas, that is, circuits of out-degree $1$, which always admit an equivalent reformulation with logarithmic depth. Hence, we construct an instance with $k\log n$ input and depth $O(\log n)$, which for $k$ large enough satisfies the requirements of statement 1.
Proof 4.2 (Proof of Theorem 4.1).
@@ 4.1 W[SAT] to Pathwidth, Treewidth, Logarithmic Tree-depth Modulators
Statement 5 of Theorem 4.1 implies statement 4.
@@
We suppose there exist constants $\epsilon>0,d>0$ and an algorithm that takes as input a Boolean formula $\psi$ (that is, a Boolean circuit of maximum out-degree $1$) with $N$ inputs and an integer $k$ and decides if $\psi$ has a satisfying assignment of weight $k$ in time $N^{(1-\varepsilon)k}|\psi|^{d}$.
@@ 4.2 Linear-depth Circuits and Constant Pathwidth Modulators
In this section we present Lemma 4.7 which, as mentioned will rely on the transformation of Boolean circuits to bounded-width branching programs given by Barrington’s theorem. Before we begin let us recall some definitions and background.
@@
Statement 2 of Theorem 4.1 implies statement 1.
@@
The proof relies heavily on Theorem 4.6: given an arbitrary Boolean circuit we will first construct an equivalent branching program and then convert this branching program into a CNF formula. The formula will have a modulator to small pathwidth because the branching program has constant width.
More precisely, suppose that for some $\varepsilon>0$ there exists an algorithm which for all $c>0$ takes as input a CNF formula $\phi$ and a modulator $M$ of its primal to pathwidth $c$, with $|M|=m$, and decides if $\phi$ is satisfiable in time $2^{(1-\varepsilon)m}|\phi|^{c^{\prime}}$, for some $c^{\prime}>0$. Let $\varepsilon^{\prime}=\frac{\varepsilon}{4c^{\prime}}$. We claim that we can obtain an algorithm which, given an $n$-input circuit $C$ of size $s$ and depth $\epsilon^{\prime}n$ decides if $C$ is satisfiable in time $2^{(1-\varepsilon^{\prime})n}s^{O(1)}$.
We first invoke Theorem 4.6 and obtain a branching program of length $O(4^{\epsilon^{\prime}n})$ and width $5$. We then construct a CNF formula whose variables are the $n$ variables $x_{1},\ldots,x_{n}$ representing the inputs to the branching program, plus the $3\ell$ variables $y_{1,t},y_{2,t},y_{3,t}$, for $t\in[\ell]$. The intuitive meaning of these variables is that the assignment to $y_{1,t},y_{2,t},y_{3,t}$ should encode the vertex which the execution of the branching program will reach in layer $t$. We make an injective mapping from the vertices of $V_{t}$ to the assignments of $y_{1,t},y_{2,t},y_{3,t}$ and add clauses to exclude assignments not associated with a vertex of $V_{t}$. We then add clauses to ensure that arcs are followed: for each assignment to $y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}$ as well as the at most five input variables used as labels for the vertices of $V_{t}$, we construct clauses ensuring the transitions conform to the arcs of the branching program. Finally, we add clauses ensuring that in layer $1$ we begin at the start vertex and in the last layer we reach that accept vertex. It is not hard to see that the constructed CNF formula is satisfiable if and only if there exists an assignment to the input variables accepted by the branching program, and hence the circuit.
We observe that in the formula we have constructed there is a modulator of size $n$ such that removing those variables results in a graph of pathwidth at most 5 (we can use a sequence of bags containing $\{y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}\}$ for all $t$). If we execute the supposed algorithm on this instance the running time will be at most $2^{(1-\varepsilon)n}|\phi|^{c^{\prime}}$. We have $|\phi|=O(4^{\epsilon^{\prime}n})=O(2^{\frac{\varepsilon n}{2c^{\prime}}})$. The algorithm therefore runs in $O(2^{(1-\frac{\varepsilon}{2})n})=O(2^{(1-\varepsilon^{\prime})n})$.
@@ 4.3 Linear-Depth Circuits and Weighted SAT
Statement 1 of Theorem 4.1 implies statement 5.
@@
Suppose there exist $\varepsilon>0,c>0$ such that satisfiability of $n$-input circuits of depth $\epsilon n$ and size $s$ can be decided in time $2^{(1-\varepsilon)n}s^{c}$.

### v2

@@ 3.2 Equivalent Formulation of the SETH for Max-SAT
Take such a component $C$ and for each assignment $s$ to its $\delta$ neighbor variables in $M$, calculate the assignment to the variables of $C$ that maximizes the number of clauses involving variables of $C$ which are satisfied (this can be done by going through the $2^{\sigma}$ possible assignments to the variables of the component). Let $w_{s}$ be the number of such clauses. Our goal is to remove the variables of $C$ from $\phi$ and add some clauses involving the variables of $N(C)$ such that for each assignment $s$ the total weight of satisfied new clauses is proportional to $w_{s}$. More precisely, we do the following: for each assignment $s$ to the variables of $N(C)$ we construct a clause falsified by this assignment (and only this assignment) to $N(C)$, and give it weight $\sum_{s^{\prime}\neq s}w_{s^{\prime}}$, that is give it weight that is equal to the sum of the weights corresponding to all other assignments to $N(C)$. We now observe that if we use the assignment $s$ for $N(C)$, the total weight we obtain from the new clauses is $w_{s}+(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$, that is, the contribution of each assignment (including $s$) is counted $2^{\delta}-2$ times, but $w_{s}$ is counted once more, as a $w_{s}$ term appears in the weight of all satisfied clauses. We increase the target value of the Max-SAT instance by $(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$ and implement the weights by taking copies of the constructed clauses. After performing this process exhaustively we are left with a Max-SAT instance with maximum arity $\max\{k,\delta\}$ and size at most $O(2^{\sigma+\delta}|\phi|)$, so executing the supposed algorithm gives the lemma.
@@ 4 Modulators to Constant Width and Linear Depth Circuits
In this section we present a sequence of reductions which show that several seemingly distinct fine-grained complexity hypotheses are actually the same. Our results are summarized in Theorem˜4.1 below, where we initially focus on the decision version of SAT (we discuss MaxSAT further below).
@@
    (LD-C-SETH is false) There is an $\varepsilon>0$ and an algorithm which takes as input a (bounded fan-in) boolean circuit of size $s$ on $n$ inputs with depth at most $\varepsilon n$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
@@
    (pwM-SETH is false) There is an $\varepsilon>0$ such that for all fixed $c>0$ we have an algorithm which takes as input a CNF formula $\phi$ and a size $m$ modulator of $\phi$ to pathwidth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
@@
    ($\log\!\textrm{td}$M-SETH is false) The same as the previous statement, but for a modulator to tree-depth at most $c\log|\phi|$.
  5. 5.
    (W[SAT]-SETH is false) There is an $\varepsilon>0$ such that there is an algorithm which takes as input a general Boolean formula $\phi$ on $n$ variables and an integer $k$ and decides if $\phi$ has a satisfying assignment that sets exactly $k$ variables to True in time $n^{(1-\varepsilon)k}|\phi|^{O(1)}$.
Among the five statements of Theorem˜4.1 three have an obvious chain of implications: solving SAT parameterized by the size of a modulator to logarithmic tree-depth is at least as hard as solving it for a modulator to constant treewidth, which is at least as hard as for a modulator to constant pathwidth. These relations follow readily from the fact that $\textrm{pw}\leq tw$ and $\textrm{td}\leq\textrm{tw}\log n$ for all graphs. What is perhaps a bit surprising is that these three parameters turn out to be of equivalent complexity.
@@
[…]
Proof 4.2 (Proof of Theorem˜4.1).
The implications $4\Rightarrow 3\Rightarrow 2$ follow immediately from the fact that on all $n$-vertex graphs we have $\textrm{tw}\leq\textrm{pw}$ and $\textrm{td}\leq\textrm{tw}\log n$. The implication $5\Rightarrow 4$ is given in Lemma˜4.3. The implication $2\Rightarrow 1$ is given in Lemma˜4.7. The implication $1\Rightarrow 5$ is given in Lemma˜4.9.
MaxSAT
In addition to the above we extend the results of Theorem˜4.1 to show equivalence also with MaxSAT parameterized by the size of a modulator to logarithmic tree-depth (hence also constant pathwidth and treewidth). We present this result in Section˜4.4. This extension is non-trivial, as it requires to generalize the basic idea used in Theorem˜4.1 (namely, eliminating variables using a tree-depth decomposition) from the realm of boolean to the realm of arithmetic circuits, because now to select the best assignment for a variable we have to calculate the number of satisfied clauses (this naturally corresponds to a circuit with $\max$ and ++ gates). The main challenge is to reduce such circuits to boolean ones without significantly increasing their depth, so that we can still obtain equivalence with the LD-C-SETH, and for this we need to rely on classical facts from circuit complexity such as the fact that computing the sum of $N$ integers on $N$ bits can be done with boolean circuits of depth only $O(\log N)$ (mentioned previously in Theorem˜2.1).
@@ 4.1 W[SAT] to Pathwidth, Treewidth, Logarithmic Tree-depth Modulators
Statement 5 of Theorem˜4.1 implies statement 4.
@@
We suppose there exist constants $\varepsilon>0,d>0$ and an algorithm that takes as input a Boolean formula $\psi$ (that is, a Boolean circuit of maximum out-degree $1$) with $N$ inputs and an integer $k$ and decides if $\psi$ has a satisfying assignment of weight $k$ in time $N^{(1-\varepsilon)k}|\psi|^{d}$.
@@ 4.2 Linear-depth Circuits and Constant Pathwidth Modulators
In this section we present Lemma˜4.7 which, as mentioned will rely on the transformation of Boolean circuits to bounded-width branching programs given by Barrington’s theorem. Before we begin let us recall some definitions and background.
@@
Statement 2 of Theorem˜4.1 implies statement 1.
@@
The proof relies heavily on Theorem˜4.6: given an arbitrary Boolean circuit we will first construct an equivalent branching program and then convert this branching program into a CNF formula. The formula will have a modulator to small pathwidth because the branching program has constant width.
More precisely, suppose that for some $\varepsilon>0$ there exists an algorithm which for all $c>0$ takes as input a CNF formula $\phi$ and a modulator $M$ of its primal graph to pathwidth $c$, with $|M|=m$, and decides if $\phi$ is satisfiable in time $2^{(1-\varepsilon)m}|\phi|^{c^{\prime}}$, for some $c^{\prime}>0$. Let $\varepsilon^{\prime}=\frac{\varepsilon}{4c^{\prime}}$. We claim that we can obtain an algorithm which, given an $n$-input circuit $C$ of size $s$ and depth $\varepsilon^{\prime}n$ decides if $C$ is satisfiable in time $2^{(1-\varepsilon^{\prime})n}s^{O(1)}$.
We first invoke Theorem˜4.6 and obtain a branching program of length $O(4^{\varepsilon^{\prime}n})$ and width $5$. We then construct a CNF formula whose variables are the $n$ variables $x_{1},\ldots,x_{n}$ representing the inputs to the branching program, plus the $3\ell$ variables $y_{1,t},y_{2,t},y_{3,t}$, for $t\in[\ell]$. The intuitive meaning of these variables is that the assignment to $y_{1,t},y_{2,t},y_{3,t}$ should encode the vertex which the execution of the branching program will reach in layer $t$. We make an injective mapping from the vertices of $V_{t}$ to the assignments of $y_{1,t},y_{2,t},y_{3,t}$ and add clauses to exclude assignments not associated with a vertex of $V_{t}$. We then add clauses to ensure that arcs are followed: for each assignment to $y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}$ as well as the at most five input variables used as labels for the vertices of $V_{t}$, we construct clauses ensuring the transitions conform to the arcs of the branching program. Finally, we add clauses ensuring that in layer $1$ we begin at the start vertex and in the last layer we reach that accept vertex. It is not hard to see that the constructed CNF formula is satisfiable if and only if there exists an assignment to the input variables accepted by the branching program, and hence the circuit.
We observe that in the formula we have constructed there is a modulator of size $n$ such that removing those variables results in a graph of pathwidth at most 5 (we can use a sequence of bags containing $\{y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}\}$ for all $t$). If we execute the supposed algorithm on this instance the running time will be at most $2^{(1-\varepsilon)n}|\phi|^{c^{\prime}}$. We have $|\phi|=O(4^{\varepsilon^{\prime}n})=O(2^{\frac{\varepsilon n}{2c^{\prime}}})$. The algorithm therefore runs in $O(2^{(1-\frac{\varepsilon}{2})n})=O(2^{(1-\varepsilon^{\prime})n})$.
@@ 4.3 Linear-Depth Circuits and Weighted SAT
Statement 1 of Theorem˜4.1 implies statement 5.
@@
Suppose there exist $\varepsilon>0,c>0$ such that satisfiability of $n$-input circuits of depth $\varepsilon n$ and size $s$ can be decided in time $2^{(1-\varepsilon)n}s^{c}$.
@@
4.4 MaxSAT and Modulators of Logarithmic Tree-depth
We are now ready to show that the results of Theorem˜4.1 remain valid if instead of SAT we consider the more general MaxSAT problem. In particular, a fast satisfiability algorithm for circuits of depth $\varepsilon n$ would be sufficient not only to decide satisfiability of formulas which are close to having logarithmic tree-depth, but also to decide the maximum number of satisfiable clauses of such formulas. This will become crucial later, because the versions of Independent Set we consider will be relatively straightforward to reduce to MaxSAT on formulas which have a small modulator to instances of constant treewidth (and hence logarithmic tree-depth).
Theorem 4.11.
The following two statements are equivalent:
  1. 1.
    (LD-C-SETH is false) There is an $\varepsilon>0$ and an algorithm which takes as input a (bounded fan-in) boolean circuit of size $s$ on $n$ inputs with depth at most $\varepsilon n$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
  2. 2.
    (MaxSAT for logarithmic tree-depth modulators) There is an $\varepsilon>0$ such that for all fixed $c>0$ we have an algorithm which takes as input a CNF formula $\phi$, a target value $t$, and a size $m$ modulator of $\phi$ to tree-depth at most $c\log|\phi|$ and decides if there exists an assignment that satisfies at least $t$ clauses of $\phi$ in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
Proof 4.12.
The implication $2\Rightarrow 1$ follows immediately from Theorem˜4.1, because an algorithm that can solve MaxSAT can also solve SAT, implying the fourth statement of Theorem˜4.1. We therefore focus on the implication $1\Rightarrow 2$. Fix an $\varepsilon>0$ for which we have an algorithm that can decide the satisfiability of boolean circuits of depth $\varepsilon n$ in time $(2-\varepsilon)^{n}s^{O(1)}$.
Fix a $c>0$ and suppose we are given a formula $\phi$ with a modulator $M$ of size $m$ such that $\phi-M$ has tree-depth at most $c\log|\phi|$. We want to find an assignment that satisfies as many clauses of $\phi$ as possible in time $(2-\varepsilon^{\prime})^{m}|\phi|^{O(1)}$, for some $\varepsilon^{\prime}>0$. We will assume without loss of generality that, for all fixed $\delta>0$ we have $|\phi|<2^{\delta m}$, as otherwise we can enumerate all assignments to $M$ in time $2^{m}<|\phi|^{1/\delta}=|\phi|^{O(1)}$ and for each such assignment compute its best extension to the rest of the instance in time $\phi^{O(1)}$ (using the fact that the remaining instance has logarithmic tree-depth). By choosing an appropriate $\delta$ we can therefore assume that any quantity that is $O(\log|\phi|)$ is upper-bounded by $\varepsilon m$.
Let us first give a high-level overview of our approach. We will begin by constructing an arithmetic circuit using bounded fan-in $\max$ and ++ gates and depth $O(\log|\phi|)$. The idea is that the gates of the circuit will allow us to eliminate all the variables of $\phi$ except those of $M$. In particular, for each variable of $\phi-M$ we have a $\max$ gate (we select the better assignment for the variable) whose inputs are ++ gates with two inputs: the number of clauses immediately satisfied by giving a specific value to the variable, and the circuit corresponding to the remaining formula. Hence, feeding an assignment to the variables of $M$ to the circuit will compute the maximum number of clauses that can be satisfied by extending this assignment. Once we have this arithmetic circuit we would like to convert it into a boolean circuit, however doing this in a naive way will increase the depth to super-logarithmic in $|\phi|$. We will therefore first “compress” the circuit: we will replace the bounded fan-in gates with gates of fan-in $\log\log|\phi|$, in the process reducing the depth to $\log|\phi|/\log\log|\phi|$. We then use the fact that the sum or maximum of $N$ numbers of $N$ bits can be computed by a boolean circuit of depth $O(\log N)$, or in our case, the sum of $\log|\phi|$ numbers on $\log|\phi|$ bits can be computed by a boolean circuit of depth $\log\log|\phi|$. Plugging this circuit in the place of high fan-in gates produces the final boolean circuit of the desired depth.
Let us proceed to the construction. We will construct a boolean circuit with $n=m$ inputs and depth at most $O(\log|\phi|)<\varepsilon m$, on which we can apply the supposed algorithm. The circuit will be satisfiable if and only if there exists an assignment satisfying $t$ clauses of $\phi$, so we will obtain the implication.
We first construct an arithmetic circuit. In such a circuit the value of each gate is an integer and we will use the operations ++ and $\max$. We describe the construction of such a circuit by induction.
Our arithmetic circuit will have as inputs the variables of $M$ with the natural correspondence between boolean and integer values, namely, False is 0 and True is $1$. In order to allow us to encode literals we will also use $\neg$ gates, which given input $x$ output 1x1-x; these will only be used in the layer immediately after the inputs, while the rest of the circuit will consist only of ++ and $\max$ gates of fan-in $2$, as well as integer constants.
We make the following inductive claim: we can produce from $\phi$ and $M$ an arithmetic circuit $C$ which satisfies the following:
  1. 1.
    For each assignment $\sigma$ to $M$, if we feed $\sigma$ to the circuit $C$ we obtain as output the number of clauses satisfied by the best assignment of $\phi$ that is consistent with $\sigma$.
  2. 2.
    If the primal graph of $\phi-M$ is connected, then the depth of $C$ is at most $3\textrm{td}(\phi-M)+2\log|\phi|+1$.
  3. 3.
    If the primal graph of $\phi-M$ is not connected, then the depth of $C$ is at most $3\textrm{td}(\phi-M)+2\log|\phi|+2$.
Claim 2.
An arithmetic circuit satisfying the above can be constructed in polynomial time, given $\phi$ and $M$ as well as a tree-depth decomposition of $\phi-M$.
{claimproof}
We show the claim by induction. For the base case, suppose $M$ contains all variables of $\phi$. Then, we can construct a circuit as follows: for each clause, construct a $\max$ gate, and feed it as input all the literals of the clause (directly or through $\neg$ gates); then add a ++ gate that computes the sum of the gates constructed for the clauses. This circuit uses ++ and $\max$ gates of unbounded fan-in, but can easily be converted to one that uses gates of fan-in $2$ and depth at most $2\log|\phi|+1$ as desired.
We now have two inductive cases, depending on whether $\phi-M$ is connected or not. First, suppose that $\phi-M$ is connected, there exists therefore a variable $x\not\in M$ such that $\textrm{td}(\phi-M-\{x\})=\textrm{td}(\phi-M)-1$. We compute two formulas $\phi_{0},\phi_{1}$, by setting $x$ to 0 or $1$ respectively and simplifying $\phi$ (that is, deleting satisfied clauses and falsified literals). Inductively construct circuits $C_{0},C_{1}$ for the two formulas. The circuit $C$ has a $\max$ gate as output; we feed into this gate two inputs, calculated by ++ gates: the first ++ gate takes as input the output of $C_{0}$ and an integer equal to the number of clauses that contain the literal $\neg x$, while the other ++ gate takes as input the output of $C_{1}$ and an integer equal to the number of clauses that contain $x$. Correctness of the new circuit can be established by induction. The depth is at most $2+\max\{d_{0},d_{1}\}$, where $d_{0},d_{1}$ are the depths of $C_{0},C_{1}$ respectively. We have $d_{0}\leq 3\textrm{td}(\phi-M-\{x\})+2\log|\phi|+2\Rightarrow 2+d_{0}\leq 3\textrm{td}(\phi-M)+2\log|\phi|+1$, as desired.
Suppose then that $\phi-M$ is disconnected and the connected components correspond to formulas $\phi_{1},\phi_{2},\ldots,\phi_{\ell}$. Intuitively, we would like to construct the circuits $C_{1},\ldots,C_{\ell}$ and then feed their output to a ++ gate, but this is slightly complicated from the fact that we want to keep the fan-in at $2$. If one of the formulas, say $\phi_{1}$ is large, that is, $|\phi_{1}|\geq|\phi|/2$, then we can proceed by building one circuit $C_{1}$ for $\phi_{1}$ and another $C_{2}$ for $\phi_{2}\cup\ldots\cup\phi_{\ell}$, and feeding their outputs to a ++ gate. The circuit $C_{1}$ has depth $1$ less than the desired bound (as $\phi_{1}$ has a connected primal graph), while $C_{2}$ has sufficiently smaller depth as its size has decreased by a factor of at least $2$, compared to $\phi$.
We are therefore left with the case where $\phi-M$ is disconnected into formulas $\phi_{1},\ldots,\phi_{\ell}$, and all these subformulas have size at most $|\phi|/2$. Partition the formulas $\phi_{1},\ldots,\phi_{\ell}$ into two sets $L,R$ by initially setting $L=\{\phi_{1},\ldots,\phi_{\ell}\}$ and $R=\emptyset$ and then repeatedly selecting an arbitrary formula from $L$ and moving it to $R$ as long as we maintain the invariant that the total size of formulas in $L$ is at least $|\phi|/2$. We now are at a situation where the total size of formulas in $R$ is at most $|\phi|/2$; $L$ is such that if we remove any formula it contains then its size becomes at most $|\phi|/2$; all formulas $\phi_{i}$ have size at most $|\phi|/2$. We now construct three circuits: a circuit $C_{R}$ representing all formulas of $R$; a circuit $C_{1}$ representing an arbitrarily chosen formula $\phi_{i}\in L$; and a circuit $C_{L}$ representing $L\setminus\{\phi_{i}\}$. We feed the outputs of $C_{1}$ and $C_{L}$ into a ++ gate, and the output of that gate with the output of $C_{R}$ into an ++ gate, which is the final output. We observe that because all three circuits represent formulas whose size is at least a factor of $2$ smaller than $\phi$, their depths are at least $3$ less than the desired bound, so the two extra layers we added allow us to stay within the target depth.
We now show how the circuit we have produced can be turned into a circuit of smaller depth by increasing the fan-in. The reason this will be helpful is that the depth of a boolean circuit required to add $N$ integers of $N$ bits is asymptotically the same as for $2$ integers of $N$ bits, so increasing the fan-in will allow us to more efficiently convert the arithmetic circuit into a boolean circuit.
Claim 3.
There is a polynomial-time algorithm that takes as input an arithmetic circuit using $\max$ and ++ gates of depth $d$ and fan-in $2$ and outputs an equivalent circuit of depth $O(d/\log d)$ using $\max$ and ++ gates of fan-in at most $O(d)$.
{claimproof}
To simplify notation we assume that $d$ is a power of $2$. Suppose without loss of generality that the given circuit is arranged in layers, where layers alternate between $\max$ and ++ gates. Suppose the layers are numbered $1,\ldots,d$ and gates at layer $i$ have inputs in layer i1i-1. We will conserve the gates in one out of each $\log d$ layers, call these layers the selected layers.
Consider now a gate $g$ at a selected layer. Its value depends on the values of at most $2^{\log d}=d$ gates in the previous selected layer. Consider then the binary tree rooted at $g$ with the $d$ gates from the previous selected layer as leaves. We can replace this tree with a constant-depth circuit of fan-in $O(d)$. In particular, our new tree will have a $\max$ gate as output, which takes as input $O(d)$ ++ gates, which each take as input some subset of the $d$ gates of the previous selected layer. This can be done inductively: for each gate $g^{\prime}$ in the path from the previous selected layer to $g$ we construct such a bounded-depth circuit: if $g^{\prime}$ is a $\max$ gate we take the circuits constructed for its inputs and merge their outputs; while if $g^{\prime}$ is a ++ gate we apply the distributive property $\max\{a,b\}+\max\{c,d\}=\max\{a+c,a+d,b+c,b+d\}$ to maintain the property that the output is a $\max$ gate. Doing this for all gates in selected layers allows us to eliminate intermediate layers and obtain the desired depth bound.
We now recall (Theorem˜2.1) that one can perform additions and maximum calculations of $N$ numbers on $N$ bits with depth $O(\log n)$ boolean circuits. Putting this together with the two claims we obtain the proof of the statement. In particular, we initially constructed a bounded fan-in arithmetic circuit of depth $O(\log|\phi|)$ and then after applying the second claim and Theorem˜2.1 converted it to an equivalent boolean circuit of depth $O(\log|\phi|)$. We add to the end a boolean comparator that checks if the output is the binary encoding of at least the target value $t$. The whole circuit now has depth $O(\log|\phi|)$, which as we explained can be assumed to be at most $\varepsilon m$, so the supposed algorithm for circuit satisfiability can be used to solve MaxSAT on the given instance.